OpenStudy (anonymous):

The St. James Theater is one of the largest theaters in New York City, with a seating capacity of 1600. They sell two different types of tickets, pre-sale and at-the-door. This combination of seats can be represented by the equation x + y = 1600. If the theater is sold out, choose the combination below that is not a possible combination of pre-sale and at-the-door tickets. Pre-sale: 1200, At-the-door: 400 Pre-sale: 522, At-the-door: 1078 Pre-sale: 847, At-the-door: 700 Pre-sale: 1600, At-the-door: 0

OpenStudy (bahrom7893):

x+y=1600 plug in each to check if it's true: 1200+400=1600 522+1078=1600 847+700=1547 <- wrong

OpenStudy (bahrom7893):

since the theater sold out, your choice is: Pre-sale: 847, At-the-door: 700

OpenStudy (anonymous):

bahrom is exactly right. Another way of thinking about the question is: Which of these answers doesn't sum to 1600? Once you translate the problem, it's easy.